# Zechen Li1,  Shohreh Deldari1,  Linyao Chen2,  Hao Xue1,  Flora D. Salim1

### v1

@@ A.6 Baselines for Task-Aware Tuning Stage
Transformer (Vaswani et al., 2017). The Transformer model is a widely-used architecture in various tasks, including time-series forecasting and classification. It uses self-attention mechanisms to capture long-range dependencies in sequential data, making it highly effective for modeling complex temporal relationships.
@@
Informer (Zhou et al., 2021). Informer is a transformer-based model designed for long sequence time-series data. It addresses key limitations of standard Transformers, such as high time complexity and memory usage, through three innovations: ProbSparse self-attention, which reduces time complexity; self-attention distilling, which enhances efficiency by focusing on dominant patterns; and a generative decoder that predicts entire sequences in a single forward pass.
@@
NS-Transformer (Liu et al., 2022). Non-stationary Transformers (NS-Transformer) tackles the issue of over-stationarization in time-series by balancing series predictability and model capability. It introduces Series Stationarization to normalize inputs and De-stationary Attention to restore intrinsic non-stationary information into temporal dependencies.
@@
PatchTST (Nie et al., 2023). PatchTST is a Transformer-based model for multivariate time series tasks, using subseries-level patches as input tokens and a channel-independent approach to reduce computation and improve efficiency. This design retains local semantics and allows for longer historical context, significantly improving long-term forecasting accuracy.
@@
TimesNet (Wu et al., 2023a). TimesNet is a versatile backbone for time series analysis that transforms 1D time series into 2D tensors to better capture intraperiod and interperiod variations. This 2D transformation allows for more efficient modeling using 2D kernels. It also introduces TimesBlock to adaptively discovers multi-periodicity and extracts temporal features from transformed 2D tensors using a parameter-efficient inception block.
@@
iTransformer (Liu et al., 2024c). iTransformer reimagines the Transformer architecture by applying attention and feed-forward networks to inverted dimensions. Time points of individual series are embedded as variate tokens, allowing the attention mechanism to capture multivariate correlations, while the feed-forward network learns nonlinear representations for each token.
@@
DeepConvLSTM (Ordóñez & Roggen, 2016). DeepConvLSTM integrates four consecutive convolutional layers followed by two LSTM layers to effectively capture both spatial and temporal dynamics in sensor data. The final output vector is passed through a fully connected layer, and the softmax function is applied to produce activity class probabilities as the model’s final output.
@@
DeepConvLSTMAttn (Murahari & Plötz, 2018). DeepConvLSTMAttn enhances the original DeepConvLSTM by integrating an attention mechanism to improve temporal modeling in HAR tasks. Instead of using the last LSTM hidden state for classification, the attention mechanism is applied to the first 7 hidden states, representing historical temporal context. These states are transformed through linear layers to generate attention scores, which are passed through softmax to produce weights. The weighted sum of the hidden states is combined with the last hidden state to form the final embedding for classification.
@@
Attend (Abedin et al., 2021). The Attend model use the latent relationships between multi-channel sensor modalities and specific activities, apply data-agnostic augmentation to regularize sensor data streams, and incorporate a classification loss criterion to minimize intra-class representation differences while maximizing inter-class separability. These innovations result in more discriminative activity representations, significantly improving HAR performance.
A.7 Evaluation Metrics for Task-Aware Tuning Stage
@@
$\text{Precision}_{i}=\frac{\text{TP}_{i}}{\text{TP}_{i}+\text{FP}_{i}}$
$\text{Recall}_{i}=\frac{\text{TP}_{i}}{\text{TP}_{i}+\text{FN}_{i}}$
@@
A.8 Sensor-Language Alignment Stage Output Examples
@@
GPT-4o’s output: 0.0s to 0.04s: consistent
0.04s to 0.06s: downward
0.06s to 0.1s: consistent
0.1s to 0.12s: downward
0.12s to 0.18s: consistent
0.18s to 0.2s: upward
0.2s to 0.24s: consistent
Number of consistent segments: 4
Number of downward segments: 2
Number of upward segments: 1

### v2

@@ A.7 Baselines for Task-Aware Tuning Stage
The Transformer model is a widely-used architecture in various tasks, including time-series forecasting and classification. It uses self-attention mechanisms to capture long-range dependencies in sequential data, making it highly effective for modeling complex temporal relationships.
Informer (Zhou et al., 2021).
@@
Informer is a transformer-based model designed for long sequence time-series data. It addresses key limitations of standard Transformers, such as high time complexity and memory usage, through three innovations: ProbSparse self-attention, which reduces time complexity; self-attention distilling, which enhances efficiency by focusing on dominant patterns; and a generative decoder that predicts entire sequences in a single forward pass.
NS-Transformer (Liu et al., 2022).
@@
Non-stationary Transformers (NS-Transformer) tackles the issue of over-stationarization in time-series by balancing series predictability and model capability. It introduces Series Stationarization to normalize inputs and De-stationary Attention to restore intrinsic non-stationary information into temporal dependencies.
PatchTST (Nie et al., 2023).
@@
PatchTST is a Transformer-based model for multivariate time series tasks, using subseries-level patches as input tokens and a channel-independent approach to reduce computation and improve efficiency. This design retains local semantics and allows for longer historical context, significantly improving long-term forecasting accuracy.
TimesNet (Wu et al., 2023a).
@@
TimesNet is a versatile backbone for time series analysis that transforms 1D time series into 2D tensors to better capture intraperiod and interperiod variations. This 2D transformation allows for more efficient modeling using 2D kernels. It also introduces TimesBlock to adaptively discovers multi-periodicity and extracts temporal features from transformed 2D tensors using a parameter-efficient inception block.
iTransformer (Liu et al., 2024c).
@@
iTransformer reimagines the Transformer architecture by applying attention and feed-forward networks to inverted dimensions. Time points of individual series are embedded as variate tokens, allowing the attention mechanism to capture multivariate correlations, while the feed-forward network learns nonlinear representations for each token.
DeepConvLSTM (Ordóñez and Roggen, 2016).
@@
DeepConvLSTM integrates four consecutive convolutional layers followed by two LSTM layers to effectively capture both spatial and temporal dynamics in sensor data. The final output vector is passed through a fully connected layer, and the softmax function is applied to produce activity class probabilities as the model’s final output.
DeepConvLSTMAttn (Murahari and Plötz, 2018).
@@
DeepConvLSTMAttn enhances the original DeepConvLSTM by integrating an attention mechanism to improve temporal modeling in HAR tasks. Instead of using the last LSTM hidden state for classification, the attention mechanism is applied to the first 7 hidden states, representing historical temporal context. These states are transformed through linear layers to generate attention scores, which are passed through softmax to produce weights. The weighted sum of the hidden states is combined with the last hidden state to form the final embedding for classification.
Attend (Abedin et al., 2021).
@@
The Attend model use the latent relationships between multi-channel sensor modalities and specific activities, apply data-agnostic augmentation to regularize sensor data streams, and incorporate a classification loss criterion to minimize intra-class representation differences while maximizing inter-class separability. These innovations result in more discriminative activity representations, significantly improving HAR performance.
Chronos+MLP.
Chronos (Ansari et al., 2024)+MLP is a baseline designed to evaluate whether the performance gains in SensorLLM are solely attributable to Chronos and the MLP. In SensorLLM, Chronos is used to generate sensor embeddings, which are then mapped by the MLP for input into the LLM to perform HAR. Since Chronos does not natively support classification tasks and only processes single-channel data, we adapt it for HAR by inputting each channel’s data separately into Chronos. The resulting sensor embeddings for all channels are then concatenated and fed into an MLP, which acts as a classifier. This setup allows us to benchmark against a simpler framework and validate the unique contributions of SensorLLM’s design.
GPT4TS (Zhou et al., 2023a).
GPT4TS is a unified framework that leverages a frozen pre-trained language model (e.g., GPT-2 (Radford et al., 2019)) to achieve state-of-the-art or comparable performance across various time-series analysis tasks, including classification, forecasting (short/long-term), imputation, anomaly detection, and few-shot/zero-sample forecasting. The authors also found that self-attention functions similarly to PCA, providing a theoretical explanation for the versatility of transformers.
A.8 Evaluation Metrics for Task-Aware Tuning Stage
@@
$\text{Precision}_{i}=\frac{\text{TP}_{i}}{\text{TP}_{i}+\text{FP}_{i}}$ (10)
$\text{Recall}_{i}=\frac{\text{TP}_{i}}{\text{TP}_{i}+\text{FN}_{i}}$ (11)
@@
A.9 Sensor-Language Alignment Stage Output Examples
@@
GPT-4o’s output: 0.0s to 0.06s: steady
0.06s to 0.12s: falling
0.12s to 0.14s: steady
0.14s to 0.16s: rising
0.16s to 0.18s: steady
Total steady segments: 3
Total rising segments: 1
Total falling segments: 1
The normalized right-lower-arm x-axis gyroscope sensor readings recorded within the 0.0 to 0.18 second timeframe are presented in this sensor data. The input data displays three individual trends, with a comprehensive change count reaching 5. The examination reveals that the data’s falling inclination endured for an aggregate of 0.06 seconds, succeeded by a rising trend for a cumulative period of 0.02 seconds, and a steady pattern for a total of 0.10 seconds. The dominant trend is steady.